OpenStudy (anonymous):

Hockey teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 53 points. They won 10 more games than they tied. How many wins and how many ties did the team have? HOw many wins did the team have? How many ties did the team have?

OpenStudy (anonymous):

2x+(x-10)=53 3x-10=53 3x=63 x=21 21 wins =42 points 21-10=11 ties=11 points 42+11=53